# Demystifying Language Model Forgetting with Low-rank Example Associations

### v7

@@ Appendix A Limitations
Nevertheless, we note that the forgetting can be defined in many possible ways, each with different focus. Some measures of forgetting, such as an aggregated number of performance drop on domain-specific downstream tasks, are often more sensible to downstream LLM users. Some existing works also separate out “shallow forgetting” of language models, where the forgetting is easily recovered by twisting the prompts (Kotha et al., 2024; Chen et al., 2024). Importantly, forgetting defined in different ways may not always align with each other. We again remind the readers of the definition of forgetting used in the paper, and we consider analysis under different definitions of forgetting to be meaningful future work.
Theoretical analysis of low-rank example associations. While we conclude that low-rank example associations are prevalent in LLM forgetting from empirical results, we do not have a theoretical guarantee on why example associations must be low rank, and in which cases the example associations become more complicated. We leave a theoretical understanding of low-rank example associations as future work. Besides, we consider that the line of research on mechanistic interpretability of LLMs (Conmy et al., 2023), which dissects subsets of neurons or computation graphs that are crucial for a prediction, can offer meaningful insights for interpreting the low-rank property.
@@
Separating out intended forgetting from unintended forgetting. Aligning with most prior works on continual learning, our experiments focus on the mitigation of forgetting. We consider the goal aligns with our experiment setup, as upstream examples from Dolma and Tulu are carefully curated and cleaned, and the newly learned tasks intend to accumulate knowledge instead of overwrite previous knowledge. However, we highlight that forgetting is not always detrimental; in fact, the models are often expected to forget (unlearn) sensitive, noisy, and outdated examples (Nguyen et al., 2022; Blanco-Justicia et al., 2025). In addition, active forgetting can serve as a strategy to enhance the plasticity of models when learning new knowledge (Chen et al., 2023). We believe that our analysis of forgetting is helpful for future work that performs targeted forgetting of knowledge.
@@ Appendix E Details of the Example Similarity Metrics
Representational similarity. We measure textual representation cosine similarity given by the (1) final layer representations of OLMo-1B, (2) all-distilroberta-v1 sentence transformer (SBERT) model (Reimers & Gurevych, 2019), and (3) OpenAI text-embedding-3-small model. We average representations of maximum 100 training examples as the representation of a task $T_{i}$.
Inner products between projected gradients and model weight updates. The increase of the log perplexity $z_{ij}$} can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee et al., 2019; Doan et al., 2020), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning. Following Park et al. (2023); Xia et al. (2024), we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson & Lindenstrauss, 1984).
Inner products between projected gradients. We also measure the negative inner products of the loss gradients between the upstream example $x_{j}$ and a learned task $T_{i}$, $z^{\textrm{g-g}}_{ij}=-\langle\nabla_{\theta}f(x_{j}),\nabla_{\theta}f(T_{i})\rangle$, as an approximation of forgetting (Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019).
@@ Appendix F Effect of Targeted Replay on New-Task Learning and Downstream Performance
LLM leaderboard task evaluation. We evaluate the change of performance on Open LLM Leaderboard tasks [^1] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about before and after fine-tuning a model. We use same set of LLM leaderboard tasks as the OLMo technical report (Groeneveld et al., 2024). For OLMo-7B models, we evaluate downstream performance of 8 models fine-tuned the Dolly tasks. For OLMo-7B Instruct, We evaluate downstream performance after fine-tuning on 4 held-out tasks from OLMo2-SFT-Mix (marked in Table 9). We compare LLM leaderboard task performance of no replay, replaying random examples, and replaying with forgetting predicted by offline MF. Table 8 summarizes results on OLMo-7B; Table 9 summarizes results on OLMo-7B-Instruct.
@@ Appendix H Effect of Model Training Setups on Forgetting
Low-rank approximations hold at various learning rates, batch sizes, and LoRA fine-tuning. Table 11 summarizes $R^{2}$ of low-rank approximations. Reducing learning rate from the default 2e-6 to 1e-6 causes very minor change of $R^{2}$; Nevertheless, we notice increased $R^{2}$ under a larger learning rate like 5e-6, indicating even simpler associations between the learned tasks and forgotten upstream examples. Under this overly large learning rate, the model suffers substantially more forgetting on all examples. Besides, increasing the batch size (under the same number of parameter update steps) causes an increase in $R^{2}$, indicating simpler associations, possibly due to reduced variance of learned models under a larger batch size. LoRA fine-tuning (Hu et al., 2022) results in simpler associations despite the model forgets less than full parameter fine-tuning.
@@ Appendix I Combining Embedding and Matrix Completion based Prediction
In this section, we examine forgetting after fine-tuning on clean, noisy, or adversarial new tasks. We experiment with Ultrafeedback (Cui et al., 2024), an instruction tuning dataset where each instruction is paired with one most preferred response and many rejected responses. For the clean variant, we fine-tune the model over the preferred responses. For the noisy variant, we shuffle the set of the preferred responses so that the instruction and the response mismatch. For the adversarial variant, we fine-tune the model over the rejected responses. We leave preference pair fine-tuning that simultaneously utilize both preferred and rejected responses as future work.
@@
We mostly apply other training setups and hyperparameters in Aljundi et al. (2019a). Each task (rotation) includes 1,000 training examples. We train a MLP classifier to predict among 10 digits given an input image without providing its rotation (or the task identifier). We experiment MLP classification models with 1 layer (a linear model) to 5 layers. We collect the example associations $Z$, and visualize them in Figure 14. The upstream and the newly learned rotation tasks are ordered by their rotations in the $x$ or $y$ axis.

### v8

@@ Appendix A Limitations
Nevertheless, we note that the forgetting can be defined in many possible ways, each with different focus. Some measures of forgetting, such as an aggregated number of performance drop on domain-specific downstream tasks, are often more sensible to downstream LLM users. Some existing works also separate out “shallow forgetting” of language models, where the forgetting is easily recovered by twisting the prompts (kotha2024understanding; chen2024jet). Importantly, forgetting defined in different ways may not always align with each other. We again remind the readers of the definition of forgetting used in the paper, and we consider analysis under different definitions of forgetting to be meaningful future work.
Theoretical analysis of low-rank example associations. While we conclude that low-rank example associations are prevalent in LLM forgetting from empirical results, we do not have a theoretical guarantee on why example associations must be low rank, and in which cases the example associations become more complicated. We leave a theoretical understanding of low-rank example associations as future work. Besides, we consider that the line of research on mechanistic interpretability of LLMs (conmy2023towards), which dissects subsets of neurons or computation graphs that are crucial for a prediction, can offer meaningful insights for interpreting the low-rank property.
@@
Separating out intended forgetting from unintended forgetting. Aligning with most prior works on continual learning, our experiments focus on the mitigation of forgetting. We consider the goal aligns with our experiment setup, as upstream examples from Dolma and Tulu are carefully curated and cleaned, and the newly learned tasks intend to accumulate knowledge instead of overwrite previous knowledge. However, we highlight that forgetting is not always detrimental; in fact, the models are often expected to forget (unlearn) sensitive, noisy, and outdated examples (nguyen2022survey; blanco2025digital). In addition, active forgetting can serve as a strategy to enhance the plasticity of models when learning new knowledge (chen2023improving). We believe that our analysis of forgetting is helpful for future work that performs targeted forgetting of knowledge.
@@ Appendix E Details of the Example Similarity Metrics
Representational similarity. We measure textual representation cosine similarity given by the (1) final layer representations of OLMo-1B, (2) all-distilroberta-v1 sentence transformer (SBERT) model (reimers-2019-sentence-bert), and (3) OpenAI text-embedding-3-small model. We average representations of maximum 100 training examples as the representation of a task $T_{i}$.
Inner products between projected gradients and model weight updates. The increase of the log perplexity $z_{ij}$} can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee2019WideNN; Doan2020ATA), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning. Following Park2023TRAKAM; Xia2024LESSSI, we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson1984ExtensionsOL).
Inner products between projected gradients. We also measure the negative inner products of the loss gradients between the upstream example $x_{j}$ and a learned task $T_{i}$, $z^{\textrm{g-g}}_{ij}=-\langle\nabla_{\theta}f(x_{j}),\nabla_{\theta}f(T_{i})\rangle$, as an approximation of forgetting (LopezPaz2017GradientEM; chaudhry2018efficient).
@@ Appendix F Effect of Targeted Replay on New-Task Learning and Downstream Performance
LLM leaderboard task evaluation. We evaluate the change of performance on Open LLM Leaderboard tasks [^2] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about before and after fine-tuning a model. We use same set of LLM leaderboard tasks as the OLMo technical report (Groeneveld2024OLMoAT). For OLMo-7B models, we evaluate downstream performance of 8 models fine-tuned the Dolly tasks. For OLMo-7B Instruct, We evaluate downstream performance after fine-tuning on 4 held-out tasks from OLMo2-SFT-Mix (marked in Table 9). We compare LLM leaderboard task performance of no replay, replaying random examples, and replaying with forgetting predicted by offline MF. Table 8 summarizes results on OLMo-7B; Table 9 summarizes results on OLMo-7B-Instruct.
@@ Appendix H Effect of Model Training Setups on Forgetting
Low-rank approximations hold at various learning rates, batch sizes, and LoRA fine-tuning. Table 11 summarizes $R^{2}$ of low-rank approximations. Reducing learning rate from the default 2e-6 to 1e-6 causes very minor change of $R^{2}$; Nevertheless, we notice increased $R^{2}$ under a larger learning rate like 5e-6, indicating even simpler associations between the learned tasks and forgotten upstream examples. Under this overly large learning rate, the model suffers substantially more forgetting on all examples. Besides, increasing the batch size (under the same number of parameter update steps) causes an increase in $R^{2}$, indicating simpler associations, possibly due to reduced variance of learned models under a larger batch size. LoRA fine-tuning (Hu2021LoRALA) results in simpler associations despite the model forgets less than full parameter fine-tuning.
@@ Appendix I Combining Embedding and Matrix Completion based Prediction
In this section, we examine forgetting after fine-tuning on clean, noisy, or adversarial new tasks. We experiment with Ultrafeedback (cuiultrafeedback), an instruction tuning dataset where each instruction is paired with one most preferred response and many rejected responses. For the clean variant, we fine-tune the model over the preferred responses. For the noisy variant, we shuffle the set of the preferred responses so that the instruction and the response mismatch. For the adversarial variant, we fine-tune the model over the rejected responses. We leave preference pair fine-tuning that simultaneously utilize both preferred and rejected responses as future work.
@@
We mostly apply other training setups and hyperparameters in Aljundi2019OnlineCL. Each task (rotation) includes 1,000 training examples. We train a MLP classifier to predict among 10 digits given an input image without providing its rotation (or the task identifier). We experiment MLP classification models with 1 layer (a linear model) to 5 layers. We collect the example associations $Z$, and visualize them in Figure 14. The upstream and the newly learned rotation tasks are ordered by their rotations in the $x$ or $y$ axis.